# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ 4.1  Evaluation Metrics for Fact-Checking Systems (RQ1)
Figure 5: Taxonomy of evaluation metrics for fact-checking systems.
Evaluating LLMs for fact-checking and related areas such as grounded generation, summarization, and error detection is a crucial and evolving field. It addresses one of the most significant challenges in LLM deployment: their tendency to “hallucinate,” or generate text that sounds plausible but is factually incorrect [10, 18]. Evaluation in this context typically involves checking the model’s outputs against provided evidence or reliable external sources [10]. Previously, a wide array of methods has been developed, including using LLMs themselves as evaluators and building benchmarks that unify datasets and tasks [10]. Figure 5 illustrates a comprehensive summary of the complete set of metrics.
@@ 4.1.1  Traditional Classification Metrics
Most commonly, the evaluation tasks are approached as classification problems, which determines whether a claim is true or identifies errors in responses [28, 29]. Metrics like Accuracy [30, 28, 31, 32, 33, 34], Precision, Recall, and F1-score [35, 21, 36, 15, 37, 38] are widely used for these tasks. In multi-class scenarios, such as classifying statements as supported, refuted, or inconclusive, macro-averaged versions of these metrics are employed [35, 21, 28]. These also serve as standard measures in detection tasks [39]. For short-form responses, token-level precision with annotated answers is typical [15].
These metrics offer quantitative performance indicators, making it easier to compare models or methods directly [21, 30, 28, 32]. They often reduce complex outputs to a binary (i.e., correct or incorrect) judgment and overlook reasoning quality or nuanced inaccuracies [18]. They may also be misleading in datasets with imbalanced labels [39].
@@ 4.1.2  Lexical and Semantic Overlap Metrics
These metrics are standard for fluency and content similarity and can capture meaning beyond exact word matches [43, 15, 42]. However, they do not measure factual correctness. High overlap or semantic scores may still correspond to factually incorrect content. Older semantic metrics may not align well with the reasoning capabilities of modern LLMs [3, 40, 41].
@@ 4.1.3  Factuality-Specific and Grounding Metrics
Specialized metrics have been developed to directly evaluate factual consistency. These go beyond surface similarity and focus on whether the model’s claims align with evidence. For example, benchmarks like LLM-AGGREFACT use detailed human annotations to assess support levels for claims [10]. ReaLMistake focuses on binary error detection, especially in reasoning and context alignment [29]. Other tools like the LEAF Fact-check Score compute the ratio of factually supported sentences to the total response [30], while Knowledge F1 (KF1) measures the overlap between human-used and model-used knowledge [15]. FactScore-Bio classifies responses based on retrieved evidence [44], and some methods aggregate multiple signals into a final factuality probability score [39]. Metrics like Insight Mastery Rate (IMR) and Justification Flaw Rate (JFR) assess explanatory quality [45]. Natural Language Inference (NLI) techniques and textual entailment tasks also serve to classify claims as supported, refuted, or unverifiable [35, 46, 47, 33, 11].
@@ 4.1.4  LLM-Based and Prompt-Based Evaluation
A growing trend involves using LLMs themselves as evaluators. This includes having LLMs classify responses as correct or flawed, and rate the factual accuracy of claims when prompted [10, 29]. The LLM-as-a-judge paradigm treats powerful language models as referees that compare and score the outputs of other models, which often produces results that closely align with human judgments [45, 50]. LLMs are also widely used in tasks such as decomposing complex claims [10, 51, 42], generating probing questions [52, 30, 51, 53], and selecting relevant evidence [52, 30, 20, 42]. While techniques like zero-shot, few-shot, Chain-of-Thought, ReAct, and HiSS are not metrics themselves, their impact is assessed using factuality metrics [21, 12, 54, 32]. Some systems even use LLMs to rate and verify retrieved documents [51], or use them to check for hallucinations [15]. The Preservation Score evaluates how much original content remains intact after hallucination correction [49]. LLMs enable more nuanced and context-sensitive evaluations than traditional metrics [55]. They can reduce human effort in evaluation tasks as well [45]. Their performance can be inconsistent due to sensitivity to prompt phrasing, and they may introduce bias or misjudgments [29, 55].
@@ 4.1.5  Human Evaluation
Despite automation advances, human evaluation remains essential, especially for complex and subjective aspects like explanation clarity and overall response quality [56]. Evaluators often use Likert scales to rate Readability, Coverage, Non-Redundancy, and Quality [21, 18, 45]. In dialogue tasks, several studies also assess aspects like Usefulness and Humanness [15]. Human-annotated data often forms the ground truth for many benchmarks [10, 29, 44], and evaluation criteria may include Redundancy, Diversity, Fairness, and Suitability [45]. Human judgments remain the gold standard, particularly for evaluating factual correctness and nuanced generation quality [29, 18], even though it is time-consuming, costly, and can introduce subjective variance depending on evaluators and criteria [29, 45].
@@ 4.1.6  Comparative Summary and Trends
The landscape of LLM evaluation is becoming increasingly sophisticated. Traditional metrics like Accuracy and F1-score still serve as foundational tools for classification tasks [35, 21, 28]. However, more advanced evaluations—focused on factuality and grounding—are gaining prominence, especially in response to challenges like hallucination [10, 30, 29, 44].
Human-annotated benchmarks and specialized metrics help ensure robustness, while LLMs are now frequently integrated into the evaluation loop—whether for scoring, verifying, or generating intermediate outputs [10, 29, 32]. Though promising, these LLM-based evaluations require careful validation against human judgments due to reliability concerns [29, 55]. Human evaluation continues to play a vital role, particularly for qualitative and high-stakes tasks. The trend is toward hybrid frameworks that combine multiple evaluation strategies (e.g., automated metrics, LLM reasoning, human oversights, etc.) to assess LLMs more holistically [20, 15, 51]. Thus, evaluating LLM fact-checking across diverse languages and modalities, including multimodal or cross-lingual fact-checking, is an emerging frontier and demands the adaptation or creation of new evaluation techniques [48, 48, 32, 57, 22].
Figure 6: Intrinsic vs. extrinsic hallucinations in LLM outputs: The source text provides verifiable ground truth about Ebola and COVID-19 vaccines. The intrinsic hallucination example contradicts the fact explicitly stated in the source, where extrinsic hallucination introduces new information not supported by the source.
@@ 4.2  Impact of Hallucinations on Fact-Checking Reliability (RQ2)
Hallucinations in LLMs refer to outputs that seem fluent, coherent, and linguistically correct but factually inaccurate, nonsensical, unsupported, or entirely fabricated [58, 59]. As these outputs are presented with the same level of confidence and linguistic fluency as factually accurate statements (see Figure 6), it is often difficult for users to detect without external verification [60, 58]. These hallucinations can occur from training corpora with contradictory, outdated, or misleading information, biases, lack of grounding, and even prompts [21, 3, 61].
@@
Table 3: An overview of studies on hallucinations covered in RQ2.
@@
Multimodal Fact-Checking 2024 Cao et al. [36], Ge et al. [43], Sharma et al. [78], Qi et al. [31], Geng et al. [57], Khaliq et al. [22] 6
Multilingual Fact-Checking 2024 Quelle et al. [48] 1
@@ 4.2.1  Hallucinations in LLMs
Nature and Types of Hallucinations. In the context of fact-checking, hallucinations manifest as intrinsic or extrinsic errors [62, 63, 64]. Intrinsic hallucinations occur when the model’s generated output contradicts the source content. In contrast, extrinsic hallucinations introduce information that cannot be verified by any provided evidence, meaning output that can neither be supported nor contradicted by the source, often fabricating details not grounded in reality [50, 60, 63].
@@
Summarization models may intrinsically hallucinate by stating a fact at odds with the source article, or an open-ended question and answer (Q&A) model may extrinsically hallucinate entirely new (false) information that it presents as factual. Authors of [3, 60, 49] demonstrated that there are generally two types (see example in Figure 7) of LLM hallucinations: (i) faithfulness, when the generated text is not faithful to the input context; and (ii) factuality, when the generated text is not factually correct for world knowledge.
@@
Causes of Hallucination. Hallucinations arise from fundamental misalignment in how LLMs are trained and used. The core training objective of most LLMs is to predict the next word in a sentence based on patterns learned from massive text data, not to guarantee truthfulness [34, 65, 66]. This means models are optimized to produce text that is coherent and contextually appropriate rather than factually accurate [15, 17]. Thus, LLMs’ tendency to hallucinate can be traced to their optimization focus on linguistic fluency and coherence, rather than factual precision, especially when faced with queries outside their training distribution or when internal knowledge conflicts arise [3]. During training, they absorb countless statements-including inaccuracies and fictional content-without an explicit mechanism to distinguish truth from false [5, 6, 67, 7, 8]. As a result, an LLM may confidently generate confident-sounding claims that align with linguistic patterns in its memory but are not grounded in facts. Wang et al. [44] demonstrate that "knowledge error" can occur when the model produces hallucinated or inaccurate information due to a lack of relevant knowledge or internalizing false knowledge in the pre-training stage or the problematic alignment process.
@@
Additionally, since an LLM’s parametric knowledge base is fixed after training, it can become outdated or insufficient, leading to guesses on topics it does not know [46, 19, 69, 70, 30]. When prompted “closed-book” (without access to external data), a model faced with an unknown fact will often fill the gap by generating a likely-sounding answer, and that is a major source of hallucination [10]. In other cases, even if grounding documents are provided, the model might improperly blend or misattribute information from those sources, causing intrinsic hallucinations [10]. Hallucinations may also be an inherent adversarial vulnerability of LLMs; even nonsense or out-of-distribution prompts can trigger the model to produce a false but fluent response [71]. This suggests that, beyond knowledge gaps, the model’s sensitivity to input perturbations and over-reliance on spurious correlations in training data can induce hallucinations.
@@
Implications for Reliability in Fact-Checking. Hallucinations directly undermine the reliability of LLMs as fact-checkers across domains, and in some cases, it can be fatal when deployed for mission-critical tasks [15, 33, 72, 49]. Studies estimate that even advanced models like GPT-4 and LLaMA-2 produce false factual statements in roughly 5–10% of their responses on general knowledge queries [34]. When LLMs are used in fact-checking systems, hallucinations can lead to the misclassification of claims, thereby underestimating the accuracy and trustworthiness of fact-checking procedures [48]. Moreover, hallucinations risk amplifying misinformation when LLM-generated content aligns with existing false narratives, unknowingly facilitating their widespread acceptance in the eyes of the public [12, 73, 53].
@@
The complexity of detecting hallucinated content, particularly when intertwined with accurate information, complicates the verification process and increases mental workload for human fact-checkers, potentially impairing decision making as evidenced by studies indicating reduced discernment following LLM-assisted fact-checking [33, 73]. Additionally, automated fact-checking systems leveraging LLMs for evidence retrieval, claim interpretation, or verdict generation are highly vulnerable, as hallucinations occurring at any stage can contaminate the entire fact-checking pipeline, ultimately resulting in unreliable outcomes [14, 3, 68]. Recent work emphasizes the need to quantify these phenomena systematically to enhance model reliability, proposing methodologies specifically aimed at measuring and addressing hallucination severity within faithfulness evaluations [50]. Singhal et al. [53] state that in terms of fact-checking, simply assigning a veracity label is inadequate; the prediction must be supported by evidence to ensure the system’s transparency and to bolster public trust.
@@ 4.2.2  Mitigation Strategies for LLM Hallucinations
Fine-tuning and Instruction Tuning. Fine-tuning pre-trained LLMs on datasets tailored to a specific domain or task, emphasizing factuality, substantially enhances their reliability [10]. Instruction tuning, a specialized variant of fine-tuning, trains models to better adhere to explicit instructions aimed at factual and verifiable responses. Specifically, domain-specific adaptation involves fine-tuning on datasets rich in factual claims and evidence from specialized areas, such as news, medical, political, religious, or legal domains, thus helping models learn the intricacies of factual language and reasoning pertinent to these fields [11]. Tang et al. [10] showed fine-tuning transformer-based models, RoBERTa, DeBERTa, and T5 models on synthetic and ANLI data boosts robustness and achieves better results. Hu et al. [12] proposed that leveraging LLMs’ ability to provide rationales could enhance finetuned small language models (SLM), ultimately improving the performance of fake news detection. Zhang et al. [28] demonstrated that in clinical claim evaluation, domain-tuned discriminative models such as BioBERT with 80.2% accuracy outperformed both zero-shot and fine-tuned generative LLMs like Llama3-70B, even after tuning. Zhao et al. [14] fine-tuned Pretrained LMs utilizing various strategies such as BERT-FC with dual loss, LIST5 with list-wise reasoning, and T5 language model, RoBERTa using NLI, and MULTIVERS with multitask learning.
@@
In terms of instruction-following techniques for factuality, several notable works, including Cheung et al. [46], explicitly enhance instruction-following models by integrating external knowledge specifically for fact-checking. They illustrated that targeted fine-tuning can significantly improve factual accuracy. Setty et al. [11] demonstrates that smaller, finely-tuned models can sometimes surpass larger, more general models in accuracy for fact-checking tasks; Tran et al. [30] introduces two parallel self-training methods to updated LLMs and boost factual reliability-Supervised Fine-Tuning (SFT) using verified responses and Simple Preference Optimization (SimPO) using fact-based ranking; Qi et al. [31] introduced a two-stage instruction tuning process which adapts InstructBLIP for OOC misinformation detection utilizing a two-stage process: aligning it to the news domain using NewsCLIPpings data [74] and fine-tuning on GPT-4-generated inconsistency explanations. Leite et al. [55] introduce a multi-stage weak supervision approach utilizing instruction-tuned LLMs prompted with 18 credibility signals to generate weak labels, which are then aggregated to predict content veracity, thereby minimizing hallucinations and enhancing transparency for human fact-checkers. Jing et al. [50] fine-tuned two HHEM DeBERTa NLI models on synthetic data and found that the HHEM model with fine-tuning on synthetic data can outperform LLMs in domain-specific evaluation, given carefully crafted training data.
@@
Retrieval-Augmented Generation (RAG). RAG has emerged as a prominent technique to ground LLM outputs in external, verifiable knowledge sources, aiming to enhance factual context, reduce hallucination, and decrease reliance on potentially flawed internal knowledge [53, 48, 3, 33]. RAG architectures generally include a retriever module that gathers relevant information from external sources, such as web documents or databases, and a generator module (the LLM) that synthesizes this retrieved information and formulates answers or assessments [15, 39].
@@
Several studies have demonstrated the effectiveness and variations of RAG. For instance, Singhal et al. [53] report how integrating external knowledge and feedback loops can significantly improve factuality; Peng et al. [15] introduced LLM-AUGMENTER that enhances LLM responses by retrieving external knowledge, linking raw evidence with related context, verifying outputs, and iteratively refining prompts using automated feedback until the response is free from hallucination and factually grounded. On the other had, Quelle et al. [48] allowed LLMs to perform Google searches to gather evidence related to the claim; Sankararaman et al. [39] presented a method Provenance where a dual stage cross-encoder framework, one to filter and weight context items and another to another to assess factuality score, aggregating these scores using weights for threshold-independent evaluation; Khaliq et al. [22] applies multi-modal reasoning within RAG proposing Chain of RAG (CoRAG) and Tree of RAG (ToRAG) specifically for political contexts;

### v2

@@ 4.1  Evaluation Metrics for Fact-Checking Systems (RQ1)
Figure 6: Taxonomy of evaluation metrics for fact-checking systems. The framework categorizes metrics into five groups. Each category is illustrated with representative measures and the number of papers adopting them, highlighting the breadth of methodologies employed in fact-checking research.
Evaluating LLMs for fact-checking and related areas such as grounded generation, summarization, and error detection is a crucial and evolving field. It addresses one of the most significant challenges in LLM deployment: their tendency to “hallucinate,” or generate text that sounds plausible but is factually incorrect [10, 18]. Evaluation in this context typically involves checking the model’s output against provided evidence or reliable external sources [10]. Previously, a wide range of methods have been developed, including using LLMs themselves as evaluators and building benchmarks that unify datasets and tasks [10]. Figure 6 illustrates a comprehensive summary of the complete set of metrics.
@@ 4.1.1  Traditional Classification Metrics
Most often, evaluation tasks are approached as classification problems, which determine whether a claim is true or identify errors in responses [28, 29]. Metrics like accuracy [28, 30, 31, 32, 33, 34], precision, recall, and F1-score [15, 21, 35, 36, 37, 38] are widely used for these tasks. In multiclass scenarios, such as classifying statements as supported, refuted, or inconclusive, macro-averaged versions of these metrics are employed [21, 28, 35]. These metrics also serve as standard measures in detection tasks [39]. For short-form responses, token-level precision with annotated answers is typical [15]. Token-level responses allow for partial evaluation, where an answer can be mostly correct, but as it is highly dependent on the specific tokenizer utilized, direct comparisons between different models are difficult.
These metrics offer quantitative performance indicators, making it easier to compare models or methods directly [21, 28, 30, 32]. They often reduce complex output to a binary (i.e., correct or incorrect) judgment, overlooking reasoning quality or nuanced inaccuracies [18]. They may also be misleading in datasets with imbalanced labels [39].
@@ 4.1.2  Lexical and Semantic Overlap Metrics
These metrics are standard for assessing fluency and content similarity, and can capture meaning beyond exact word matches [15, 42, 43]. However, they do not measure the factual correctness. High overlap or semantic scores may still correspond to factually incorrect content. Older semantic metrics may not align well with the reasoning capabilities of modern LLMs [3, 40, 41].
@@ 4.1.3  Factuality-Specific and Grounding Metrics
Specialized metrics have been developed to directly evaluate the factual consistency. These go beyond surface similarity and focus on whether the model’s claims align with the evidence. For example, benchmarks like LLM-AGGREFACT use detailed human annotations to assess support levels for claims [10]. It is a meta-benchmark for factuality that aggregates 11 different publicly available fact-checking and hallucination datasets. ReaLMistake focuses on binary error detection, especially in reasoning and context alignment [29]. Other tools like the LEAF fact-check Score compute the ratio of factually supported sentences to the total response [30] by decomposing each response into individual sentences, then each sentence is independently verified against retrieved external knowledge sources. Knowledge F1 (KF1), utilized by Peng et al. [15], measures the overlap between human-used and model-used knowledge. FactScore-Bio classifies responses based on retrieved evidence [44], and some methods aggregate multiple signals into a final factuality probability score [39]. Metrics such as Insight Mastery Rate (IMR) and Justification Flaw Rate (JFR) assess explanatory quality [45] where IMR represents the proportion of low-scoring fact-checking responses relative to the total number of questions, where a Grade of three or below (on a ten-point scale) indicates errors in the target LLM’s response. On the other hand, JFR denotes the percentage of cases where the target LLM conducted correct verdict prediction yet had poor justification, based on the conditions set by IMR. Natural Language Inference (NLI) techniques and textual entailment tasks also serve to classify claims as supported, refuted, or unverifiable [11, 33, 35, 46, 47].
@@ 4.1.4  LLM-Based and Prompt-Based Evaluation
A growing trend involves using LLMs themselves as evaluators. This includes having LLMs classify responses as correct or flawed, and rate the factual accuracy of claims when prompted [10, 29]. The LLM-as-a-judge paradigm treats powerful language models as referees that compare and score the outputs of other models, which often produces results that closely align with human judgments [45, 50]. LLMs are also widely used for tasks such as decomposing complex claims [10, 42, 51], generating probing questions [30, 51, 52, 53], and selecting relevant evidence [20, 30, 42, 52]. While techniques like zero-shot, few-shot, Chain-of-Thought, ReAct, and HiSS are not metrics themselves, their impact is assessed using factuality metrics [12, 21, 32, 54]. Some systems even use LLMs to rate and verify retrieved documents [51], or use them to check for hallucinations [15]. The Preservation Score evaluates how much original content remains intact after hallucination correction [49]. LLMs enable more nuanced and context-sensitive evaluations than traditional metrics [55]. They can reduce human effort in evaluation tasks as well [45]. Their performance can be inconsistent due to sensitivity to prompt phrasing, and they may introduce bias or misjudgments [29, 55].
@@ 4.1.5  Human Evaluation
Despite automation advances, human evaluation remains essential, especially for complex and subjective aspects like explanation clarity and overall response quality [56]. Evaluators often use Likert scales to rate Readability, Coverage, Non-Redundancy, and Quality [18, 21, 45]. In dialogue tasks, several studies also assess aspects like Usefulness and Humanness [15]. Human-annotated data often forms the ground truth for many benchmarks [10, 29, 44], and evaluation criteria may include Redundancy, Diversity, Fairness, and Suitability [45]. Human judgments remain the gold standard, particularly for evaluating factual correctness and nuanced generation quality [18, 29], even though it is time-consuming, costly, and can introduce subjective variance depending on the evaluators and the criteria used [29, 45].
@@ 4.1.6  Comparative Summary and Trends
The landscape of LLM evaluation is becoming increasingly sophisticated. Traditional metrics like Accuracy and F1-score still serve as foundational tools for classification tasks [35, 21, 28]. However, more advanced evaluations focused on factuality and grounding are gaining prominence, especially in response to challenges such as hallucinations [10, 29, 30, 44].
Human-annotated benchmarks and specialized metrics help ensure robustness, while LLMs are now frequently integrated into the evaluation loop, whether to score, verify, or generate intermediate outputs [10, 29, 32]. Although promising, these LLM-based evaluations require careful validation against human judgments due to reliability concerns [29, 55]. Human evaluation continues to play a vital role, particularly in high-stakes and qualitative tasks. The trend is toward hybrid frameworks that combine multiple evaluation strategies (e.g., automated metrics, LLM reasoning, and human oversight) to assess LLMs more holistically [15, 20, 51]. Thus, evaluating LLM fact-checking across diverse languages and modalities, including multimodal or cross-lingual fact-checking, is an emerging frontier and demands the adaptation or creation of new evaluation techniques[22, 32, 48, 57].
Figure 7: Intrinsic vs. extrinsic hallucinations in LLM outputs: The source text provides verifiable ground truth about Ebola and COVID-19 vaccines. The intrinsic hallucination example contradicts the fact explicitly stated in the source, whereas extrinsic hallucination introduces new information that is not supported by the source.
@@ 4.2  Impact of Hallucinations on Fact-Checking Reliability (RQ2)
Hallucinations in LLM refer to results that seem fluent, coherent, and linguistically correct but are factually inaccurate, nonsensical, unsupported, or completely fabricated [58, 59]. As these outputs are presented with the same level of confidence and linguistic fluency as factually accurate statements (see Figure 7), it is often difficult for users to detect them without external verification [58, 60]. These hallucinations can arise from contradictory, outdated, or misleading information in training corpora, as well as biases, lack of grounding, and even from prompts [3, 21, 61].
@@
Table 3: An overview of studies on hallucinations covered in RQ2. The table organizes prior research by types and causes of hallucinations, their implications for reliability in fact-checking, and mitigation strategies. For each category, the publication years, representative authors, and study counts are reported.
@@
Multimodal Fact-Checking 2024 Cao et al. [36], Ge et al. [43], Yao et al. [77], Papadopoulos et al. [78], Sharma et al. [79], Qi et al. [31], Wang et al. [80], Kakizaki et al. [81], Geng et al. [57], Khaliq et al. [22] 10
Multilingual Fact-Checking 2024 Quelle et al. [48], Zhang et al. [82], Shafayat et al. [83], Siino et al. [84], Shcharbakova et al. [85], Jannah et al. [86] 6
@@ 4.2.1  Hallucinations in LLMs
Nature and Types of Hallucinations. In the context of fact-checking, hallucinations manifest as intrinsic or extrinsic errors [62, 63, 64]. Intrinsic hallucinations occur when the model’s generated output contradicts the source content. In contrast, extrinsic hallucinations introduce information that cannot be verified by any provided evidence, resulting in output unsupported by the source and often containing fabricated details [50, 60, 63].
@@
Summarization models may intrinsically hallucinate by stating a fact that contradicts the source article, or an open-ended question and answer (Q&A) model may extrinsically hallucinate entirely new (false) information that it presents as factual. The authors of [3, 49, 60] demonstrated that LLM hallucinations can generally be classified into two types (see example in Figure 8): (i) faithfulness, where the generated text does not accurately reflect the input context; and (ii) factuality, where the generated text is factually incorrect according to real-world knowledge.
@@
Causes of Hallucination. Hallucinations arise from fundamental misalignment in the way LLMs are trained and used. The core training objective of most LLMs is to predict the next word in a sentence based on patterns learned from massive text data, not to guarantee truthfulness [34, 65, 66]. This means that models are optimized to produce text that is coherent and contextually appropriate rather than factually accurate [15, 17]. Thus, LLMs’ tendency to hallucinate can be traced to their optimization focus on linguistic fluency and coherence, rather than factual precision, especially when faced with queries outside of their training distribution or when internal knowledge conflicts arise [3]. During training, they absorb a large number of statements, including inaccuracies and fictional content, without an explicit mechanism to distinguish truth from falsehood [5, 6, 7, 8, 67]. As a result, an LLM may generate claims that sound confident and align with linguistic patterns in its memory, but are not grounded in facts. Wang et al. [44] demonstrate that a "knowledge error" can occur when the model produces hallucinated or inaccurate information due to a lack of relevant knowledge or internalizing false knowledge in the pretraining stage or the problematic alignment process.
@@
Furthermore, since the LLM’s parametric knowledge base is fixed after training, it can become outdated or insufficient, leading to guesses on topics that it does not know [19, 30, 46, 68, 69]. When prompted “closed-book” (without access to external data), a model faced with an unknown fact will often fill the gap by generating a plausible answer, which is a major source of hallucination [10]. In other cases, even if grounding documents are provided, the model might improperly blend or misattribute information from these sources, causing intrinsic hallucinations [10]. Hallucinations may also be an inherent adversarial vulnerability of LLMs. Even nonsense or out-of-distribution prompts can trigger the model to produce a false but fluent response [70]. This suggests that, beyond knowledge gaps, the model’s sensitivity to input perturbations and over-reliance on spurious correlations in training data can induce hallucinations.
@@
Implications for Reliability in Fact-Checking. Hallucinations directly undermine the reliability of LLMs as fact-checkers across domains, and in some cases, they can be fatal when deployed for mission-critical tasks [15, 33, 49, 71]. Studies estimate that even advanced models like GPT-4 and LLaMA-2 produce inaccurate factual statements in approximately 5–10% of their responses to general knowledge queries [34]. When LLMs are used in fact-checking systems, hallucinations can lead to misclassification of claims, which can lead to an underestimation of the accuracy and trustworthiness of fact-checking procedures [48]. Moreover, hallucinations risk amplifying misinformation when LLM-generated content aligns with existing false narratives, unknowingly facilitating their widespread acceptance in the eyes of the public [12, 53, 72].
@@
The complexity of detecting hallucinated content, particularly when intertwined with accurate information, complicates the verification process and increases the mental workload for human fact checkers, potentially affecting decision-making, as evidenced by studies that indicate reduced discernment after LLM-assisted fact-checking [33, 72]. Additionally, automated fact-checking systems comprising LLMs for evidence retrieval, claim interpretation, or verdict generation are highly vulnerable, as hallucinations at any stage can contaminate the entire fact-checking pipeline, ultimately resulting in unreliable outcomes [3, 14, 44]. Recent work emphasizes the need to quantify these phenomena systematically to enhance model reliability, proposing methodologies specifically aimed at measuring and addressing hallucination severity within faithfulness evaluations [50]. Singhal et al. [53] state that, in terms of fact-checking, simply assigning a veracity label is inadequate. The prediction must be supported by evidence to ensure the system’s transparency and to bolster public trust.
@@ 4.2.2  Mitigation Strategies for LLM Hallucinations
Fine-tuning and Instruction Tuning. Fine-tuning pre-trained LLMs on datasets tailored to a specific domain or task, emphasizing factuality, substantially enhances their reliability [10]. Quantitative comparisons supporting this observation are summarized in Table 7. Table 4 illustrates different fact-checking frameworks and their demonstrated approach to mitigate hallucination. Instruction tuning, a specialized variant of fine-tuning, trains models to better adhere to explicit instructions aimed at factual and verifiable responses. Specifically, domain-specific adaptation involves fine-tuning on datasets rich in factual claims and evidence from specialized areas, such as news, medical, political, religious, or legal domains, thus helping models learn the intricacies of factual language and reasoning pertinent to these fields [11]. Tang et al. [10] showed that fine-tuning transformer-based models, RoBERTa, DeBERTa, and T5 models on synthetic and ANLI data boosts robustness and improves performance. Hu et al. [12] proposed that utilizing LLMs’ ability to provide rationales could enhance fine-tuned small language models (SLMs), thereby improving fake news detection performance. Zhang et al. [28] demonstrated that in clinical claim evaluation, domain-tuned discriminative models such as BioBERT with 80.2% accuracy outperformed both zero-shot and fine-tuned generative LLMs like Llama3-70B, even after tuning. Zhao et al. [14] fine-tuned Pretrained LMs utilizing various strategies such as BERT-FC with dual loss, LIST5 with list-wise reasoning, and T5 language model, RoBERTa using NLI, and MULTIVERS with multitask learning.
Table 4: An overview of notable fact-checking frameworks for LLMs and their approach to mitigate hallucination.
Framework Hallucination Types Addressed
MiniCheck Specifically targets factual errors where LLM-generated text is not grounded in or is inconsistent with provided source documents. It is designed to handle complex errors that require synthesizing information across multiple sentences and verifying multiple facts within a single claim.
CliniFact Addresses factual errors in the highly specialized medical domain by verifying clinical research claims. It also targets errors in Logical Reasoning, as it evaluates logical statements derived from hypothesis testing in clinical trials.
VisualFactChecker Explicitly designed to mitigate hallucinations in image captioning, such as generating descriptions of non-existent objects (content) or incorrect attributes (shape, color), by using visual tools (object detection, VQA) for verification.
FACT-GPT Addresses the challenge of recurring misinformation by matching new social media posts to previously debunked claims. It tackles variations of false claims that entail or contradict a known falsehood.
Hierarchical Step-by-Step (HiSS) This method uses a prompting technique to mitigate hallucination, where prompts are given to LLMs to do fine-grained checking of news claims. Here, prompts are given to the LLM to decompose the claim into subclaims and verify each subclaim step-by-step by raising and answering a series of questions. For each question, a prompt is again given to the LLM to assess if it is confident to answer it or not, and if not, the question is given to a web search engine. The search results are then inserted back into the ongoing prompt to continue the verification process.
LLM-KG Framework Mitigates hallucinations by grounding LLM responses in a structured and verifiable Knowledge Graph. It helps prevent the generation of plausible but incorrect facts by providing explicit, verified triplets as context.
LLM-AUGMENTER Mitigates hallucinations by augmenting a fixed LLM with external knowledge to generate grounded responses. It uses automated feedback to iteratively revise responses and improve their factuality score against the evidence.
OpenFactCheck Evaluates and identifies various types of factual errors, including Knowledge Errors (hallucinated/inaccurate information), Over-commitment Errors (failing to recognize false premises), and Disability Errors (outdated information).
MEDICO This is a framework specifically designed for hallucination detection and correction. It fuses evidence from multiple sources (web, knowledge bases, knowledge graphs) to detect factual errors, provide a rationale, and iteratively revise the hallucinated content.
Yours Truly This framework is designed to fact-check social media claims by breaking down compound sentences into atomic claims and verifying each against a real-time database of fact-checked articles. It addresses the challenge of hallucinations in LLMs by verifying outputs against an external, curated knowledge source (FactStore).
Logical Consistency Framework This framework addresses inconsistencies in LLM responses, which are a vulnerability related to hallucination. It assesses and improves the logical consistency of LLMs when performing fact-checking against Knowledge Graphs on complex queries involving logical operators, aiming to make the LLM less likely to hallucinate by ensuring its reasoning is logically sound.
@@
In terms of instruction-following techniques for factuality, several notable works, including Cheung et al. [46], explicitly enhanced instruction-following models by integrating external knowledge specifically for fact-checking. They illustrated that targeted fine-tuning can significantly improve factual accuracy. Setty et al. [11] demonstrated that smaller, finely-tuned models can sometimes surpass larger, more general models in accuracy for fact-checking tasks. Tran et al. [30] introduced two parallel self-training methods to update LLMs and boost factual reliability: Supervised Fine-Tuning (SFT) using verified responses, and Simple Preference Optimization (SimPO) using fact-based ranking. Qi et al. [31] proposed a two-stage instruction tuning method to adapt InstructBLIP for OOC misinformation detection, first aligning it to the news domain using NewsCLIPpings data [73] and then fine-tuning it on GPT-4-generated inconsistency explanations. Leite et al. [55] introduced a multi-stage weak supervision approach utilizing instruction-tuned LLMs prompted with 18 credibility signals to generate weak labels, which are then aggregated to predict content veracity, thereby minimizing hallucinations and enhancing transparency for human fact checkers. Jing et al. [50] fine-tuned two HHEM DeBERTa NLI models on synthetic data and found that the HHEM model with fine-tuning on synthetic data can outperform LLMs in domain-specific evaluation, given carefully crafted training data.
@@
Retrieval-Augmented Generation. RAG has emerged as a prominent technique to ground LLM outputs in external, verifiable knowledge sources, with the aim of improving the factual context, reducing hallucination, and reliance on potentially flawed internal knowledge [3, 33, 48, 53, 89]. RAG architectures generally include a retriever module that gathers relevant information from external sources, such as web documents or databases, and a generator module (the LLM) that synthesizes this retrieved information and formulates answers or assessments [15, 39].
@@
Several studies have demonstrated the effectiveness and variations of RAG. For instance, Singhal et al. [53] reported how integrating external knowledge and feedback loops can significantly improve factuality. Table 7 illustrates these performance gains across multiple benchmarks, including PolitiFact and PubMedQA. Peng et al. [15] introduced LLM-AUGMENTER that enhances LLM responses by retrieving external knowledge, linking raw evidence with related context, verifying outputs, and iteratively refining prompts using automated feedback until the response is free from hallucination and factually grounded. On the other hand, Quelle et al. [48] allowed LLMs to perform Google searches to gather evidence related to the claim. Sankararaman et al. [39] presented a method, Provenance, where a dual-stage cross-encoder framework, one stage filters and weights context items, and the other assesses the factuality score, with the scores aggregated for threshold-independent evaluation. Khaliq et al. [22] applied multimodal reasoning within RAG, proposing Chain of RAG (CoRAG) and Tree of RAG (ToRAG), specifically for political contexts.